OpenStudy (anonymous):

find the derivative of the function by using the definition of the derivative f(x)=2/x

OpenStudy (anonymous):

Start by using:\[\lim_{h \rightarrow 0} [2/(x + h) - 2/x] / h\]

OpenStudy (anonymous):

Then, for the expression between the brackets, use (x + h)x as a common denominator. The "h" will cancel out.

OpenStudy (anonymous):

so it would be 2(x+h)-2(x)/h

OpenStudy (anonymous):

and it would be 4x^2

OpenStudy (anonymous):

Between the brackets, and just looking at that portion for now, 2/(x+h) - 2/x = [2x - 2(x+h)] / [x(x+h)]

OpenStudy (anonymous):

So, first simplify (after seeing how I got that equation) and plug that into the limit expression inside the brackets.

OpenStudy (anonymous):

so I'd put 0 for it

OpenStudy (anonymous):

Are you saying that 2x - 2(x+h) is 0?

OpenStudy (anonymous):

yeah…?

OpenStudy (anonymous):

2x - 2(x+h) = 2x - 2x - 2h. And that does not equal 0. What does it equal?

OpenStudy (anonymous):

-2?

OpenStudy (anonymous):

2x - 2(x+h) = 2x - 2x - 2h = (2x - 2x) - 2h = 0 - 2h. What does that equal?

OpenStudy (anonymous):

0-2=-2

OpenStudy (anonymous):

You are disregarding the "h". Look again. I did not write 0 - 2. I am asking you what 0 - 2h equals.

OpenStudy (anonymous):

-2h???? ahhh im confused

OpenStudy (anonymous):

It's OK to be confused, as long as you can "see the light at the end of the tunnel". Many sources of math confusion are just little hurdles that, once over, you never stumble on again. Just think of "h" as another variable, similar to "x" or "y", that's all. And yes, you are making progress. In the solution to this problem, this intermediate answer is -2h so far. But we're not done yet.

OpenStudy (anonymous):

So, the "between the brackets" expression, 2/(x+h) - 2/x = [2x - 2(x+h)] / [x(x+h)] now equals -2h / [x(x+h)]. So now, you can put that into the limit expression in my first post.

OpenStudy (anonymous):

So, you should have\[\lim_{h \rightarrow 0} ([-2h] / [x(x + h)]) / h\]

OpenStudy (anonymous):

I dont understand where the their 2 came from

OpenStudy (anonymous):

third*

OpenStudy (anonymous):

This simplifies to lim h→0 (−2h)/[x(x+h)h]. So, divide out the factor of "h" now (NOT the "x+h").

OpenStudy (anonymous):

nvm I goyt it

OpenStudy (anonymous):

Okay, good so far. After you divide out the "h" factor, you've got lim h→0 (−2)/[x(x+h)]. So, now the h can go to zero without giving "0" in the denominator.

OpenStudy (anonymous):

The denominator of x(x+h) will become x(x+0) or x(x) which is x^2

OpenStudy (anonymous):

And since the numerator is just -2, you put numerator and denominator together for your answer and you're done!

OpenStudy (anonymous):

h is replace by 0 so it disappears?

OpenStudy (anonymous):

"h" disappears only at the end when you take the limit. "h" stays as "h" up until the last step (taking the limit) all while you are doing the arithmetic and the simplifying. For most of the problem, "h" is not 0 and is just a variable.

OpenStudy (anonymous):

A good way to conceptualize "h" is "change in x" or "delta x" or :\[\Delta x\]

OpenStudy (anonymous):

right. ok I think I got this now. Thankyou so much for your help, and for taking the time to explain this tome . YOu were a huge help. thanks again..

OpenStudy (anonymous):

Starting to make a little sense? I could give you a little more mathematical theory if that would help. With some students, that helps. Others just want to know the "how to get the answer" part.

OpenStudy (anonymous):

how about finding the slope of A tangent line..?

OpenStudy (anonymous):

Once you know how to get the derivative that we just did, the slope of a tangent line is a real snap. It's just the derivative! And if you need a concrete number at a certain value of x, you just plug that value into the x and you have the slope "at that point". If you want the slope in general, then just leave the "x" as "x".

OpenStudy (anonymous):

The derivative for a function is itself just another function. And you could take the derivative of that derivative function. It would end up being the second derivative of the original function.